From GlueXWiki
Jump to: navigation, search

October 27, 2011 FDC meeting


  1. Production Construction Tracking
    • Status (Dave)
    • First package tests (Lubomir)
  2. Engineering update (?)
  3. Electronics update (Chris)
  4. Testing set-up at EEL126 (Beni)
  5. TDR (Lubomir)
  6. Other


Participants: Eugene, Mark, Dave, Chris, Vladimir, Simon, and Lubomir.


- Dave: Work on second package. Started stringing second wire frame. Chris will put the elements on the first wire plane tomorrow followed by wire deadening. Now gluing the daughter cards on the 3th and 4th cathodes. Anatoly made a board connected to a box that makes it easier to measure the resistance of the strips-daughter board connection. Such measurement is now a standard procedure to check the gluing of the rig-flexes. Problems with the Parker controller: the ADC connected to the laser sensor gives voltages in the range 0-2V instead of 0-10V. Dave will discuss with Larry possible options to fix the problem.

- Lubomir: Fighting with the noise on the top cell of the first package. Before we identified two HV sectors that induce noise on one top and one bottom connector. The noise is induced on the traces leading to the connector, not on the strips itself. The top cathode was first rotated by 15 deg w.r.t. standard position to pin down the point where the noise originates. In the suspected areas three resistors ("big" 1MOhm resistors at the HV side) were replaced, but this didn't help. This morning we rotated the top cathode by 30deg in which case many traces leading to one of the connectors cover the suspected area. In this configuration we were able to pin down the source of the noise with higher precision. For both sectors it is at place where the +HV trace crosses the solid part of the G10 on the wire frame. This is on the back of the boards and after lamination we have no access to this place. So, most likely there will be no fix unless we can rotate the cathodes by 180deg so that the connectors are covering the signal boards, not the HV ones. Most likely the placement of the connectors in this case will create problems, that will be discussed with Bill.


- Discussing without Bill the protective covers: aluminum 1/8" thick. Eugene asked why they are so thick so we can't use iron source. Lubomir: we didn't require the possibility to use a source, but actually when testing the package it will be very easy to remove the top cover and do tests with a source. Dave had a concern that the covers are too heavy and may cut the plastic rods, but it turned out that there will be plastic inserts that will cover the rods.


- Chris: set-up is ready to start testing the first 22 pre-amps manually. It will take several days but Chris will not be at JLab next week. We will look if some of our techs, or Vladimir can do the tests. Chris wanted to know if we can have also additional brackets for the spare pre-amps.

Chamber testing at 126

- Beni: need to make 36 twisted pair ribbon cables to be connected between the translator boards and the F1TDCs. Chris will look for the cables and connectors. The production of the translator boards is in the procurement. We should check on this later.


- There's "Tracking" section before the CDC and FDC detectors that Simon will look into it. However, mainly we will concentrate on the FDC description as a detector. After the meeting Lubomir, Simon, and Eugene discussed the FDC section in more details.